OpenStudy (anonymous):

simplify; Log3 9x^2-Log3 7xy= Log3 9x/7y???

OpenStudy (turingtest):

\[\log_{3} (9x^2)-\log_{3}(7xy)=\log_{3}(9x/7y)\] ???? is the expression?

OpenStudy (anonymous):

no the term after the = is my answer

myininaya (myininaya):

it looks good to me

OpenStudy (turingtest):

agreed

OpenStudy (anonymous):

really???? wow

OpenStudy (anonymous):

how did u get this answer?

OpenStudy (anonymous):

i used the logarithm rule for subtraction

OpenStudy (anonymous):

9x^2/7xy =9x/7y

OpenStudy (anonymous):

logx-logy=log(x/y)

OpenStudy (anonymous):

this is the subtraction rule

OpenStudy (turingtest):

logarithmic properties: \[\log(ab)=\log(a)+\log(b)\] and \[\log(x^n)=n \log(x)\]

OpenStudy (anonymous):

where log is with base 3

OpenStudy (turingtest):

just make sure the bases of the logs are the same, or else those rules won't work.

OpenStudy (anonymous):

so was i wrong???

OpenStudy (turingtest):

No! you were right, I don't know what uzma did.